What is a better buy? A square pizza measuring 8 inches by 8 inches that cost $10 or a round pizza with a 9-inch diameter that also cost $10?
1
Expert's answer
2012-08-10T09:41:35-0400
The area of a square pizza is

As = 8*8 = 64 square inches

and area of a round pizza is

Ar ≈ 3.14*(9/2)² = 63.6173 square inches.

So, Ar<As and it's better to buy square pizza.